Shift report 20030318 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Mar 18 2003 - 16:05:54 EST

  • Next message: Flemming Videbaek: "Wednesday access."
    ------------------------------------------------------------
    Date: Tue Mar 18 16:05:53 2003, Shift: 08:00-16:00
    Supervisor: sjs, with: 
    Log book#14 pages 67 - 68
    
    Shift Summary:
    ------------------------------------------------------------
    Started shift with run 8487.  This is a continuation of
    run 8486 that had some settings wrong.  Both runs are 
    1/8B field, but without BFS (HV crate failure).  After
    store BRAHMS was given access to fix the crate.  Around
    noon we were ready to run with BBS, but RHIC has been
    having trouble getting a new beam store.  The FS is
    now set at 8 det, 1/4B field.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Mar 18 2003 - 16:06:52 EST